1) An economy has a saving function (s) which is given by the equation s= -300 + 0.2Y. If the level of national income is 3000 what is the APS?


2) If, in an economy, out of every additional 100 of national income 20 is taxed. 20 is spent and10 is spent on imports, the value of the multiplier is?


3) National income =1000 and injection = 75 calculate the new national income if the

i) moc = 0.4 and mpt =0.1

li) mpc = 0.2 and mpt = 0.2



4) A man has a rise in income of 23000 to 26000. As a result his saving rises from 800 to 1200. His mpc is?



1
Expert's answer
2021-11-28T17:56:26-0500

Solution:

1.). APS is calculated by dividing total savings by income level. Normally, disposable (after-tax) income is used.

S = -300 + 0.2Y = -300 + 0.2(3,000) = -300 + 600 = 300

Savings = 300

Income = 3000

APS = "\frac{300}{3000} = 0.1"


APS = 0.1

 

2.). First, determine MPC:

MPC is equal to the percentage of new income spent on consumption

For every additional $100, $20 is spent, which is the consumption.

MPC = "\frac{20}{100} = 0.2"

 

Multiplier = "\frac{1}{(1 - MPC)} = \frac{1}{(1 - 0.2)} = \frac{1}{0.8} = 1.25"


Multiplier = 1.25

 

3.). i.). The new national income = Multiplier x Change in AD

Multiplier = "\frac{1}{(1 - MPC)} = \frac{1}{(1 - 0.4)} = \frac{1}{0.6} = 1.67"


The new national income = 1.67 "\times" 1075 = 1,795.25

The new national income = 1,795.25

 

ii.). The new national income = Multiplier x Change in AD

Multiplier = "\frac{1}{(1 - MPC)} = \frac{1}{(1 - 0.2)} = \frac{1}{0.8} = 1.25"

The new national income = 1.25 "\times" 1075 = 1,343.75

The new national income = 1,343.75

 

4.). First, calculate the marginal propensity to save, which is equal to ΔS "\div" ΔY

MPS = "\frac{(1,200 - 800)}{(26,000 - 23,000)} = \frac{400}{3,000} = 0.13"


MPC = 1 – MPS = 1 – 0.13 = 0.87


MPC = 0.87